# From Rankings to Insights: Evaluation Should Shift Focus from Leaderboard to Feedback

### v2

@@ C.2.1 Seed Data
Prompt of Quality Annotation (reasoning) ## Task
Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label.
## Rule
Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Problem-Solving**: Does it require devising a solution or strategy?
6. **Deep Thinking**: Does it require deep reasoning and consideration of multiple factors?
7. **Novelty:** Does the question present a unique or unusual scenario that the LLM is unlikely to have encountered before?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
@@ C.2.2 Synthesize New Queries
Prompt of RealMix I will provide three real user questions and a reference question’s type tags. Your task is to create a new question in the {domain} domain with the same tags as reference question. To ensure the question feels authentic, you should utilize real-world details drawn from the three real user questions.
### You MUST follow these steps to generate the new question:
#### Task1: Real-world Details Selection:
Select a few real-world details (such as the real people, objects, scenes, settings, and any other details mentioned) from three real user questions that fit the given tags. If there are no suitable details, return [[I cannot generate a question based on the provided real user questions.]] and stop.
#### Task2: New Question Generation:
1. Although you should use details from real user questions, you must not mention the real user question in the new question.
2. The new question should be complex and challenging, requiring deep understanding and analysis of the subject. The length of the question should be at least as long as the reference question but should not be overly simplistic or repetitive. The question should be singular, not a multi-task question.
3. The new question must be **completely self-contained**, so that others can answer it without any additional information.
4. Analyze how to create the new question with chosen real-world details and provided tags. While multiple tags are available, the newly generated question only needs to align with some of them, not all. Even if the original question already fits, generate a different version.
### Output Format:
[Anylysis]: You should first complete the anylysis of task1 and task2 here.
[Question]: Summarize the newly generated question in the following format: <new_query>Insert the final new question here.</new_query>
<|begin_of_reference_query|>
{reference_query}
<|end_of_reference_query|>
<|begin_of_reference_query_type_tags|>
{type_tags}
<|end_of_reference_query_type_tags|>
<|begin_of_real_user_query_1|>
{query1}
<|end_of_real_user_query_1|>
<|begin_of_real_user_query_2|>
{query2}
<|end_of_real_user_query_2|>
<|begin_of_real_user_query_3|>
{query3}
<|end_of_real_user_query_3|>
@@ C.3 Detailed Description of Visualization and Analysis Toolkits
Analysis. We have implemented four analysis tools, with more under development.
  • Automatic Weakness Reporter presents a list of notable strengths and weaknesses detected in each model.
  • Automatic Weakness Analyzer generates natural language summaries of a model’s strengths and weaknesses.
  • Failure Mode Explorer identifies whether a model’s failure in a subtask is systematic (low ranking variance among sibling nodes) or sporadic (high variance).
  • Automatic Difference Reporter compares multiple models and highlights significant anomalies between their performances in a structured list.
C.4 Feedbacker-D-V0
@@ C.4.1 Query Quality: Human Evaluation
Table 6: Query quality and reality evaluation results by human rater.
@@ C.4.2 Query Quality: Data Contamination
Following Auto-Arena [19], we employ two methods to detect test data contamination. The first is the string match metric, as used in GPT-4 [71], which flags a data point as contaminated if any of three randomly sampled 50-character substrings (or the entire string, if shorter) from the evaluation sample appears as a substring in the training set. The second method, inspired by Platypus [72], uses sentence embedding similarity: a question is considered contaminated if its BERT-Large [73] embedding has a cosine similarity above 0.8 with any training item, enabling detection of paraphrased overlaps.
@@
C.4.3 Examples of Feedbacker-D-V0
We present several samples of Feedbacker-D-V0 in Tab. 7.
Table 7: Examples from different domains in Feedbacker-D-V0.
@@ C.4.3 Examples of Feedbacker-D-V0
C.4.4 Feedbacker-HPD
@@
Although the seed data includes human preference labels, prior work [15] has raised concerns about their reliability due to variability in human judgment. To enhance label accuracy, we enlisted two additional graduate-level annotators and determined the final preference label via majority voting, thereby mitigating individual annotator bias. Our annotation interface is illustrated in Fig. 32.
@@
The distribution of Feedbacker-HPD is shown in Fig. 33. Similar to Feedbacker-D-V0, these queries span six domains: writing, roleplay, knowledge, coding, mathematics, and reasoning.
@@
Figure 33: Domain distribution of Feedbacker-HPD.
@@ Appendix D Evaluation Methods
**Explanation:**
- **Correctness:** Assistant A’s answer directly addresses the user’s requirement to track the *number of open tabs* on the same domain using ‘localStorage‘. It provides a straightforward ‘TabTracker‘ class with methods to increment/decrement the count and store it in ‘localStorage‘. While it lacks automatic triggering of decrement on tab closure (relying on manual calls in the example), it correctly structures the core logic for counting tabs. The problem does not explicitly require automatic event handling, so the code is technically correct for the stated goal.
- **Helpfulness:** The code is concise and focused on the core requirement. The explanation clearly describes how each method works and how to use the class. The example usage demonstrates the basic workflow, making it easy to integrate into a project.
- **Assistant B’s Issues:** While B’s approach is more detailed (tracking individual tabs with IDs and URLs), it diverges from the user’s request to track *open tabs as a count*. B’s implementation requires managing individual tabs and resolving complex edge cases (e.g., identifying which tab closed), which complicates the task. Additionally, B’s ‘_handleTabClose‘ method admits it cannot reliably remove tabs, making the solution incomplete for the stated problem.
- **Key Flaw in B’s Answer:** The user asked for a *count* of open tabs, not detailed tracking of individual tabs. B’s answer overcomplicates the solution by focusing on tab metadata rather than the numeric count.
**Conclusion:** Assistant A’s answer is better because it aligns directly with the user’s requirement, provides a functional core, and is simpler and more relevant. Assistant B’s answer, while thorough, solves a different problem (tracking individual tabs) and introduces unnecessary complexity.
@@
**Explanation:**
Both assistants provided valid HTML documents with step-by-step instructions for jump-starting a car. However, **Assistant B’s answer is better** for the following reasons:
1. **Self-Contained Diagram:**
- **Assistant B** uses an embedded SVG diagram directly in the HTML, fulfilling the user’s requirement for an "all in one document." The SVG illustrates the connection process without relying on an external image file.
- **Assistant A** references an external image (‘car_battery_jumpstart.png‘), which requires the user to provide or replace the image path. This makes the document incomplete as-is.
2. **Step-by-Step Clarity:**
- **Assistant B** breaks down the steps into a more granular 8-step process, explicitly separating each cable connection into distinct steps (e.g., connecting each end of the red and black cables separately). This enhances clarity for users unfamiliar with the process.
- **Assistant A** combines some steps (e.g., all cable connections under step 3 with sub-steps), which is concise but slightly less explicit.
3. **Safety and Correctness:**
- Both answers include correct procedural steps (e.g., grounding the negative cable on the dead car’s metal surface instead of the terminal).
- **Assistant B** adds a final step (step 8) explicitly detailing the order to remove cables, which is important for safety.
4. **HTML Structure:**
- **Assistant B’s** use of SVG ensures the diagram is part of the HTML document, whereas **Assistant A’s** reliance on an external image breaks the "all in one" requirement.
**Verdict:** While both answers are correct, **Assistant B’s** self-contained implementation of the diagram and clearer step breakdown make it the superior choice.

### v3

@@ C.2.1 Seed Data
Prompt of Quality Annotation (reasoning) ## Task Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label. ## Rule Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization? 4. **Real-World Application**: Is the question something that would be encountered in real-world? 5. **Problem-Solving**: Does it require devising a solution or strategy? 6. **Deep Thinking**: Does it require deep reasoning and consideration of multiple factors? 7. **Novelty:** Does the question present a unique or unusual scenario that the LLM is unlikely to have encountered before? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>
@@ C.2.2 Synthesize New Queries
Prompt of RealMix I will provide three real user questions and a reference question’s type tags. Your task is to create a new question in the {domain} domain with the same tags as reference question. To ensure the question feels authentic, you should utilize real-world details drawn from the three real user questions. ### You MUST follow these steps to generate the new question: #### Task1: Real-world Details Selection: Select a few real-world details (such as the real people, objects, scenes, settings, and any other details mentioned) from three real user questions that fit the given tags. If there are no suitable details, return [[I cannot generate a question based on the provided real user questions.]] and stop. #### Task2: New Question Generation: 1. Although you should use details from real user questions, you must not mention the real user question in the new question. 2. The new question should be complex and challenging, requiring deep understanding and analysis of the subject. The length of the question should be at least as long as the reference question but should not be overly simplistic or repetitive. The question should be singular, not a multi-task question. 3. The new question must be **completely self-contained**, so that others can answer it without any additional information. 4. Analyze how to create the new question with chosen real-world details and provided tags. While multiple tags are available, the newly generated question only needs to align with some of them, not all. Even if the original question already fits, generate a different version. ### Output Format: [Anylysis]: You should first complete the anylysis of task1 and task2 here. [Question]: Summarize the newly generated question in the following format: <new_query>Insert the final new question here.</new_query> <|begin_of_reference_query|> {reference_query} <|end_of_reference_query|> <|begin_of_reference_query_type_tags|> {type_tags} <|end_of_reference_query_type_tags|> <|begin_of_real_user_query_1|> {query1} <|end_of_real_user_query_1|> <|begin_of_real_user_query_2|> {query2} <|end_of_real_user_query_2|> <|begin_of_real_user_query_3|> {query3} <|end_of_real_user_query_3|>
@@ C.3 Detailed Description of Visualization and Analysis Toolkits
Analysis. We have implemented four analysis tools including Automatic Weakness Reporter, Automatic Weakness Analyzer, Failure Mode Explorer, and Automatic Difference Reporter.
(1) Automatic Weakness Reporter detects capability imbalances by comparing a model’s ranking at a specific evaluation node with its overall ranking. If the local ranking differs from the global ranking by more than a predefined threshold (set to 5 in our experiments), the node is flagged as a weakness (negative deviation) or as a strength (positive deviation). This enables rapid localization of domains where the model exhibits pronounced strengths or weaknesses. For example, Gemma-3-4B-IT shows its largest deviation in root.mathematics.Task Types.Visualization.Geometric Drawing, where its overall ranking is 9 but its ranking in Geometric Drawing drops to 16, revealing a substantial weakness in spatial and geometric visualization tasks.
(2) Automatic Weakness Analyzer builds on the results of the Automatic Weakness Reporter and uses an advanced LLM (QwQ-32B in our study) to transform numerical deviations into human-readable summaries, grouping related deficiencies and strengths. This allows evaluators to understand the qualitative nature of the weaknesses detected. For Gemma-3-4B-IT, for instance, the analyzer output includes: “Coding & Tool Usage: Struggles with Swift, Batch, Lua programming languages; IDE configuration and test development.” Such summaries enable model developers to target specific technical capabilities for improvement.
(3) Failure Mode Explorer investigates whether underperformance in a hierarchical category is comprehensive or localized. For any underperforming parent node, it calculates the standard deviation (STD) of rankings across its child nodes. If the STD is within the top 20% of all nodes, the performance is considered unstable, meaning that low performance may be caused by a few sub-capabilities. If the STD is in the bottom 20%, the weakness is comprehensive and spread across all sub-capabilities. For example, gemma ranks 7 places worse in root.mathematics.Mathematical Subfields.Algebra than its overall ranking, with a low STD of 0.4899. This indicates broad underperformance across Algebra, suggesting the need for holistic improvement rather than fixes in isolated subfields such as Linear Algebra.
(4) Automatic Difference Reporter enables model-to-model comparison by measuring differences in performance variability across evaluation nodes. The larger the difference in standard deviation between two models, the more divergent their capabilities are in that domain. This is useful for identifying areas of specialization or distinct model behavior. For instance, a comparison between claude-3.7-sonnet and gpt-4o reveals substantial differences in root.roleplay.Style Types.Experimental Style.Glitch Art, root.roleplay.Theme Types.Emotional.Erotic, and root.reasoning.Application Domains.Personal Development Guidance, pointing to contrasting strengths and specialization patterns between the two models.
Overall, this integrated framework transforms large-scale ranking data into actionable insights. By combining deviation detection, natural language summarization, hierarchical stability analysis, and comparative difference measurement, it supports interpretable evaluation and targeted optimization for LLM development.
C.4 SCAN-D-V0
@@ C.4.1 Query Quality: Human Evaluation
Table 9: Query quality and reality evaluation results by human rater.
@@ C.4.2 Query Quality: Data Contamination
Following Auto-Arena Zhao et al. (2024), we employ two methods to detect test data contamination. The first is the string match metric, as used in GPT-4 Achiam et al. (2023), which flags a data point as contaminated if any of three randomly sampled 50-character substrings (or the entire string, if shorter) from the evaluation sample appears as a substring in the training set. The second method, inspired by Platypus Lee et al. (2023), uses sentence embedding similarity: a question is considered contaminated if its BERT-Large Devlin et al. (2019) embedding has a cosine similarity above 0.8 with any training item, enabling detection of paraphrased overlaps.
@@
C.4.3 Examples of SCAN-D-V0
We present several samples of SCAN-D-V0 in Tab. 10.
@@ C.4.3 Examples of SCAN-D-V0
Table 10: Examples from different domains in SCAN-D-V0.
C.4.4 SCAN-HPD
@@
Although the seed data includes human preference labels, prior work Zheng et al. (2023) has raised concerns about their reliability due to variability in human judgment. To enhance label accuracy, we enlisted two additional graduate-level annotators and determined the final preference label via majority voting, thereby mitigating individual annotator bias. Our annotation interface is illustrated in Fig. 30.
@@
The distribution of SCAN-HPD is shown in Fig. 31. Similar to SCAN-D-V0, these queries span six domains: writing, roleplay, knowledge, coding, mathematics, and reasoning.
@@
Figure 31: Domain distribution of SCAN-HPD.
@@ Appendix D Evaluation Methods
**Explanation:** - **Correctness:** Assistant A’s answer directly addresses the user’s requirement to track the *number of open tabs* on the same domain using ‘localStorage‘. It provides a straightforward ‘TabTracker‘ class with methods to increment/decrement the count and store it in ‘localStorage‘. While it lacks automatic triggering of decrement on tab closure (relying on manual calls in the example), it correctly structures the core logic for counting tabs. The problem does not explicitly require automatic event handling, so the code is technically correct for the stated goal. - **Helpfulness:** The code is concise and focused on the core requirement. The explanation clearly describes how each method works and how to use the class. The example usage demonstrates the basic workflow, making it easy to integrate into a project. - **Assistant B’s Issues:** While B’s approach is more detailed (tracking individual tabs with IDs and URLs), it diverges from the user’s request to track *open tabs as a count*. B’s implementation requires managing individual tabs and resolving complex edge cases (e.g., identifying which tab closed), which complicates the task. Additionally, B’s ‘_handleTabClose‘ method admits it cannot reliably remove tabs, making the solution incomplete for the stated problem. - **Key Flaw in B’s Answer:** The user asked for a *count* of open tabs, not detailed tracking of individual tabs. B’s answer overcomplicates the solution by focusing on tab metadata rather than the numeric count. **Conclusion:** Assistant A’s answer is better because it aligns directly with the user’s requirement, provides a functional core, and is simpler and more relevant. Assistant B’s answer, while thorough, solves a different problem (tracking individual tabs) and introduces unnecessary complexity.
@@
**Explanation:** Both assistants provided valid HTML documents with step-by-step instructions for jump-starting a car. However, **Assistant B’s answer is better** for the following reasons: 1. **Self-Contained Diagram:** - **Assistant B** uses an embedded SVG diagram directly in the HTML, fulfilling the user’s requirement for an "all in one document." The SVG illustrates the connection process without relying on an external image file. - **Assistant A** references an external image (‘car_battery_jumpstart.png‘), which requires the user to provide or replace the image path. This makes the document incomplete as-is. 2. **Step-by-Step Clarity:** - **Assistant B** breaks down the steps into a more granular 8-step process, explicitly separating each cable connection into distinct steps (e.g., connecting each end of the red and black cables separately). This enhances clarity for users unfamiliar with the process. - **Assistant A** combines some steps (e.g., all cable connections under step 3 with sub-steps), which is concise but slightly less explicit. 3. **Safety and Correctness:** - Both answers include correct procedural steps (e.g., grounding the negative cable on the dead car’s metal surface instead of the terminal). - **Assistant B** adds a final step (step 8) explicitly detailing the order to remove cables, which is important for safety. 4. **HTML Structure:** - **Assistant B’s** use of SVG ensures the diagram is part of the HTML document, whereas **Assistant A’s** reliance on an external image breaks the "all in one" requirement. **Verdict:** While both answers are correct, **Assistant B’s** self-contained implementation of the diagram and clearer step breakdown make it the superior choice.